A 2kg block is pulled with uniform speed up a plane which makes 30° with the horizontal force of 15 N which is parallel to the plane. Find the coefficient of kinetic friction between the block and the plane


1
Expert's answer
2021-07-08T06:23:01-0400

"µ_k = \frac{f_k}{ N}\\
µ_k= \frac{mg sin α − ma}{ mg cos α}\\
µ_k= - tan α − \frac{a}{ g cos α}\\
µ_k= -tan α − \frac{\frac{F}{ m}}{ g cos α}\\
µ_k= - tan30 − \frac{\frac{15}{ 2}}{ 9.8 cos 30}\\
µ_k= - -0.306349\\
µ_k= 0.306349"